# Homological algebra of pro-Lie Polish abelian groups

### v1

@@ 3.5. Injective and projective pro-Lie Polish abelian groups
Recall that an abelian Lie group $A$ is of the form $V\oplus T\oplus D$ where $V$ is a finite-dimensional vector group, $T$ is a finite-dimensional torus, and $D$ is discrete. (This is equivalent to the assertion that $A$ is a locally compact Polish abelian group with no small subgroups; see [29, Theorem 2.4].) It easily follows that there exists a surjective homomorphism $\mathbb{R}^{n}\oplus\mathbb{Z}^{\left(\omega\right)}\rightarrow A$ for some $n\in\omega$. Recall that the projective objects in $\mathbf{LiePAb}$ are precisely those of the form $V\oplus F$ where $V$ is a finite-dimensional vector group and $F$ is a countable free abelian group [29, Theorem 3.3]. The injective objects in $\mathbf{LiePAb}$ are precisely those of the form $V\oplus T$ where $V$ is a finite-dimensional vector group and $T$ is a finite-dimensional torus [29, Theorem 3.2].
@@
Suppose that $\mathcal{A}$ is a countably complete quasi-abelian category, and let $\mathcal{D}$ be a class of projective objects in $\mathcal{A}$ closed under direct sums. Suppose that $\left(G_{n}\right)$ is a tower of objects of $\mathcal{A}$ with admissible epimorphisms $\pi_{n+1}:G_{n+1}\rightarrow G_{n}$ as bonding maps. We also set $G_{-1}=0$ and $\pi_{0}=0$. Suppose that for every $n\in\omega$ there exists a admissible epimorphism $P\rightarrow\mathrm{\mathrm{Ker}}\left(\pi_{n}:G_{n}\rightarrow G_{n-1}\right)$ for some $P\in\mathcal{D}$. Define $G=\left(G_{n}\right)_{n\in\omega}\in\mathrm{pro}_{\omega}\left(\mathcal{A}\right)$. Then there exists an admissible epimorphism $D\rightarrow G$ where $D=\left(D_{n}\right)\in\mathrm{pro}_{\omega}\left(\mathcal{A}\right)$ is a tower with $D_{n}\in\mathcal{D}$ and admissible epimorphisms $p_{n+1}:D_{n+1}\rightarrow D_{n}$ as bonding maps.
@@
We define by recursion objects $D_{n}$ of $\mathcal{D}$, admissible epimorphisms $p_{n+1}:D_{n+1}\rightarrow D_{n}$, and $\varphi_{n}:D_{n}\rightarrow G_{n}$ such that $\pi_{n+1}\varphi_{n+1}=\varphi_{n}p_{n+1}$. We have that $\varphi_{0}$ exists by hypothesis. Suppose that $\varphi_{i}$ and $p_{i}$ have been defined for $i\leq n$. Then we consider a pushout diagram
@@
Since $\mathcal{A}$ is quasi-abelian, we have that $Y_{n+1}\rightarrow D_{n}$ and $Y_{n+1}\rightarrow G_{n}$ are admissible epics. Since $D_{n}$ is projective, we have that $Y_{n+1}\cong D_{n}\oplus\mathrm{\mathrm{Ker}}\left(\pi_{n+1}\right)$. By the inductive hypothesis, and since $\mathcal{D}$ is closed under direct sums, there exists an admissible epimorphism $D_{n+1}\rightarrow Y_{n+1}$. This concludes the recursive construction. ∎
@@
Since $\mathbf{proLiePAb}$ is equivalent to $\mathrm{pro}_{\omega}\left(\mathbf{LieAb}\right)$ by Proposition 3.12, by Lemma 3.39 it suffices to prove that the conclusion holds when $G$ is an abelian Lie group, in which case the conclusion follows from the above remarks. ∎
@@
Suppose that $\left(C_{i}\right)_{i\in\omega}$ is a sequence of countable abelian groups, $C=\prod_{i\in\omega}C_{i}$, and $A$ is a countable abelian Polish group. If $\mathrm{Ext}_{{}_{\mathrm{Yon}}}\left(C_{i},A\right)=0$ for every $i\in\omega$, then $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)=0$.
@@
$\mathrm{Ext}_{\mathrm{Yon}}\left(C,A_{0}\right)\rightarrow\mathrm{Ext}_{\mathrm{Yon}}\left(C^{>n},A\right)\text{.}$
@@
Suppose that $c:C\times C\rightarrow A$ is a continuous cocycle. By hypothesis, there exists a continuous function $\varphi_{0}:C\rightarrow A$ such that $c_{0}:=\delta\varphi_{0}+c$ defines an element of $\mathrm{Ext}_{\mathrm{c}}\left(C,V_{1}\right)$. Proceeding in this fashion, we can define a sequence $c_{n}\in\mathrm{Ext}_{{}_{\mathrm{Yon}}}\left(C,V_{n}\right)$ and continuous functions $\varphi_{n}:C^{\omega}\rightarrow V_{n}$ with $c_{0}=c$ such that $c_{n+1}=\delta\varphi_{n}+c_{n}$ for every $n\in\omega$.
@@
Suppose that $\left(C_{i}\right)_{i\in\omega}$ is a sequence of countable free abelian groups. Then, setting $C:=\prod_{k\in\omega}C_{k}$, we have that $C$ is projective for non-Archimedean pro-Lie abelian Polish groups.
@@
(2$\Rightarrow$(1) We have that $\mathbb{R}^{\omega}$ is projective by Lemma 3.29. We have that $\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\omega}$ is projective by Lemma 3.46. Any group of the form $\mathbb{Z}^{\alpha}\oplus\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\beta}\oplus\mathbb{R}^{\gamma}$ for $\alpha,\beta,\gamma\in\omega+1$ is a direct summand of $\mathbb{R}^{\omega}\oplus\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\omega}$, and hence projective.
@@
(3$\Rightarrow$(1) Let $\left(N_{k}\right)$ be a cofinal sequence in $\mathcal{N}\left(G\right)$ such that $G^{\left(k\right)}:=G/N_{k}$ is projective in $\mathbf{LieAb}$ for every $k\in\omega$. Then by Lemma 3.34 we have that $G_{0}=G_{\mathbb{R}}$ and $G\left/G_{0}\right.\cong\mathrm{\mathrm{lim}}_{k}G_{\mathbb{Z}}^{\left(k\right)}$ where $(G_{\mathbb{Z}}^{\left(k\right)})_{k\in\omega}$ is an epimorphic tower of free abelian groups. Whence, $G_{\mathbb{Z}}$ is isomorphic to a product of free abelian groups and hence projective in $\mathbf{proLiePAb}$. Hence, $G\cong G_{\mathbb{R}}\oplus G_{\mathbb{Z}}$ is projective in $\mathbf{proLiePAb}$.
@@
We have that $\mathbf{proLiePAb}$ has enough projectives by Theorem 3.47 and 3.40. We now prove that closed subgroups of projectives are projective. Suppose that $G$ is a closed subgroup of $\left(F\oplus\mathbb{R}\right)^{\omega}$ where $F$ is a countable free abelian group of infinite rank. Then we have that $G_{\mathbb{S}^{1}}=0$ and $G_{\mathrm{t}}=0$. Hence, we have that $G\cong G_{\mathbb{R}}\oplus G_{\mathbb{Z}}$. Since $G_{\mathbb{R}}$ is injective, we can assume without loss of generality that $G=G_{\mathbb{Z}}$.
@@
We consider the cohomological derived functor of $\mathrm{H}^{0}\circ\mathrm{Hom}^{\bullet}$ for locally compact Polish abelian groups introduced in [5]; see also [21, Section 4]. For every locally compact Polish abelian group $A$, we have an exact sequence
@@
Thus, we have that $D=D\cap\mathbb{T}^{\omega}\subseteq\mathbb{T}^{\omega}$. Thus, we have that the inclusion $D\rightarrow\mathbb{T}^{\omega}$ induces a surjective homomorphism $\mathbb{Z}^{\omega}\rightarrow D^{\vee}$. Since $D^{\vee}$ is compact and connected, it is uncountable whenever it is nonzero. Hence, $D^{\vee}=0$ and $D=0$. ∎
@@ 4.1. Pro-$p$ Polish abelian groups
Let us say that a pro-$p$ Polish group is a pro-Lie Polish abelian topological $p$-group. We let $\mathbf{PAb}\left(p\right)$ be the category of pro-$p$ Polish abelian groups, which is a thick subcategory of $\mathbf{proLiePAb}$ by Lemma 2.4. A locally compact Polish abelian group $G$ is pro-$p$ if and only if it has a basis of zero neighborhoods consisting of subgroups $U$ such that $G/U$ is a $p$-group. A pro-Lie Polish abelian group is a pro-$p$ group if and only if $G/N$ is pro-$p$ for every $N\in\mathcal{N}\left(G\right)$.
@@
For an ordinal $\alpha$, as a particular instance of [32, Theorem 1.5] in the case of the cotorsion functor with enough projectives $G\mapsto p^{\alpha}G$, we have the following:
@@
due to the fact that $G/H$ is countable, it suffices to prove that, for every $k\in\omega$, the image of $[c]$ in $\mathrm{Ext}(\left(\mathbb{Z}/p\mathbb{Z}\right)^{\mathbb{N}},H)$ induced by the inclusion $\mathbb{Z}/p\mathbb{Z}\rightarrow\left(\mathbb{Z}/p\mathbb{Z}\right)^{\mathbb{N}}$ in the $k$-th coordinate, is nonzero. As this image is the element $[c_{k}]$ represented by the cocycle $c_{k}$, we have that it is nontrivial by the choice of $c_{k}$. This shows that $\mathrm{Ext}(\left(\mathbb{Z}/p\mathbb{Z}\right)^{\mathbb{N}},G)\neq 0$.
@@
(1$\Rightarrow$(2) By Theorem 4.2, it suffices to prove that $G$ is divisible has a basis of zero neighborhoods consisting of divisible subgroups. As in the proof of (1$\Rightarrow$(2) of Theorem 4.2, we have that $G$ is divisible. Suppose that $G$ does not have a basis of zero neighborhoods consisting of divisible subgroups. Then it has a basis $\left(V_{k}\right)_{k\in\omega}$ of compact open subgroups that are not divisible and such that $d\left(V_{k}\right)$ is not open. For $k\in\omega$, let $\rho_{k}\geq 1$ be the least countable ordinal such that $p^{\rho_{k}}V_{k}$ is not open. By compactness of $V_{k}$, we have that $\rho_{k}\leq\omega$. Indeed, if $\rho_{k}>\omega$, then we have that $\left\{p^{n}V\setminus p^{n+1}V:n\in\omega\right\}$ is an infinite clopen partition of the compact set $V\setminus\rho^{\omega}V$, which is impossible. Thus, proceeding as in the proof of (1$\Rightarrow$(2) of Theorem 4.2, one obtains that there exists a sequence $\left(T_{k}\right)$ of finite $p$-groups such that, setting $T:=\prod_{k\in\omega}T_{k}$, one has that $\mathrm{Ext}\left(T,G\right)\neq 0$, which is a contradiction.
@@
Suppose that $A$ is an injective pro-$p$ Polish abelian group, $B$ is a Polish abelian group, and $\pi:A\rightarrow B$ is a surjective continuous homomorphism. Since $A$ is divisible, $B$ is also divisible. If $\left(V_{k}\right)$ is a basis of zero neighborhoods of $A$ consisting of divisible subgroups, then $\left(\pi\left(V_{k}\right)\right)$ is a basis of zero neighborhoods of $B$ consisting of divisible subgroups. Thus, $B$ is an injective pro-$p$ Polish abelian group. This shows that quotients of injectives are injective in $\mathbf{PAb}\left(p\right)$, and hence $\mathbf{PAb}\left(p\right)$ has homological dimension at most $1$. ∎

### v2

@@ 3.5. Injective and projective pro-Lie Polish abelian groups
Recall that an abelian Lie group $A$ is of the form $V\oplus T\oplus D$ where $V$ is a finite-dimensional vector group, $T$ is a finite-dimensional torus, and $D$ is discrete. (This is equivalent to the assertion that $A$ is a locally compact Polish abelian group with no small subgroups; see [30, Theorem 2.4].) It easily follows that there exists a surjective homomorphism $\mathbb{R}^{n}\oplus\mathbb{Z}^{\left(\omega\right)}\rightarrow A$ for some $n\in\omega$. Recall that the projective objects in $\mathbf{LiePAb}$ are precisely those of the form $V\oplus F$ where $V$ is a finite-dimensional vector group and $F$ is a countable free abelian group [30, Theorem 3.3]. The injective objects in $\mathbf{LiePAb}$ are precisely those of the form $V\oplus T$ where $V$ is a finite-dimensional vector group and $T$ is a finite-dimensional torus [30, Theorem 3.2].
@@
Suppose that $\mathcal{A}$ be a quasi-abelian subcategory of $\mathbf{PAb}$ closed under countable products, and let $\mathcal{D}$ be a class of projective objects in $\mathcal{A}$ closed under direct sums. Suppose that $\left(G_{n}\right)$ is an inverse sequence of objects of $\mathcal{A}$ with surjective continuous homomorphisms $\pi_{n+1}:G_{n+1}\rightarrow G_{n}$ as bonding maps. We also set $G_{-1}=0$ and $\pi_{0}=0$. Suppose that for every $n\in\omega$ there exists a surjective continuous homomorphism $P\rightarrow\mathrm{\mathrm{Ker}}\left(\pi_{n}:G_{n}\rightarrow G_{n-1}\right)$ for some $P\in\mathcal{D}$. Then there exists a surjective continuous homomorphism $\mathrm{lim}_{n}D_{n}\rightarrow\mathrm{lim}_{n}G_{n}$ where $D=\left(D_{n}\right)$ is an inverse sequence with $D_{n}\in\mathcal{D}$ and surjective continuous homomorphisms $p_{n+1}:D_{n+1}\rightarrow D_{n}$ as bonding maps.
@@
We define by recursion objects $D_{n}$ of $\mathcal{D}$, continuous homomorphisms $p_{n+1}:D_{n+1}\rightarrow D_{n}$, and $\varphi_{n}:D_{n}\rightarrow G_{n}$ such that $\pi_{n+1}\varphi_{n+1}=\varphi_{n}p_{n+1}$. We have that $\varphi_{0}$ exists by hypothesis. Suppose that $\varphi_{i}$ and $p_{i}$ have been defined for $i\leq n$. Then we consider a pushout diagram
@@
Since $D_{n}\rightarrow G_{n}$ and $G_{n+1}\rightarrow G_{n}$ are continuous surjective homomorphisms, the same holds for $Y_{n+1}\rightarrow D_{n}$ and $Y_{n+1}\rightarrow G_{n}$. Since $D_{n}$ is projective, we have that $Y_{n+1}\cong D_{n}\oplus\mathrm{\mathrm{Ker}}\left(\pi_{n+1}\right)$. By the inductive hypothesis, and since $\mathcal{D}$ is closed under direct sums, there exists a surjective continuous homomorphism $D_{n+1}\rightarrow Y_{n+1}$. This concludes the recursive construction.
One then has that $\mathrm{lim}_{n}\varphi_{n}:\mathrm{lim}_{n}D_{n}\rightarrow\mathrm{lim}_{n}G_{n}$ is a continuous surjective homomorphism. ∎
@@
By Lemma 3.43 it suffices to prove that the conclusion holds when $G$ is an abelian Lie group, in which case the conclusion follows from the above remarks. ∎
@@
Suppose that $\left(C_{i}\right)_{i\in\omega}$ is a sequence of countable abelian groups, $C=\prod_{i\in\omega}C_{i}$, and $A$ is a countable abelian group. If $\mathrm{Ext}_{{}_{\mathrm{Yon}}}\left(C_{i},A\right)=0$ for every $i\in\omega$, then $\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)=0$.
@@
$\mathrm{Ext}_{\mathrm{Yon}}\left(C,A\right)\rightarrow\mathrm{Ext}_{\mathrm{Yon}}\left(C^{>n},A\right)\text{.}$
@@
Suppose that $c:C\times C\rightarrow A$ is a continuous cocycle. By hypothesis, there exists a continuous function $\varphi_{0}:C\rightarrow A$ such that $c_{0}:=\delta\varphi_{0}+c$ defines an element of $\mathrm{Ext}_{\mathrm{c}}\left(C,V_{1}\right)$. Proceeding in this fashion, we can define a sequence of continuous $2$-cocycles $c_{n}:C\times C\rightarrow V_{n}$ and continuous functions $\varphi_{n}:C^{\omega}\rightarrow V_{n}$ with $c_{0}=c$ such that $c_{n+1}=\delta\varphi_{n}+c_{n}$ for every $n\in\omega$.
@@
Suppose that $\left(C_{i}\right)_{i\in\omega}$ is a sequence of countable free abelian groups. Then, setting $C:=\prod_{k\in\omega}C_{k}$, we have that $C$ is projective for pro-Lie Polish abelian groups.
@@
(2$\Rightarrow$(1) We have that $\mathbb{R}^{\omega}$ is projective by Lemma LABEL:Lemma:R-projective. We have that $\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\omega}$ is projective by Lemma 3.50. Any group of the form $\mathbb{Z}^{\alpha}\oplus\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\beta}\oplus\mathbb{R}^{\gamma}$ for $\alpha,\beta,\gamma\in\omega+1$ is a direct summand of $\mathbb{R}^{\omega}\oplus\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\omega}$, and hence projective.
@@
(3$\Rightarrow$(1) Let $\left(N_{k}\right)$ be a cofinal sequence in $\mathcal{N}\left(G\right)$ such that $G^{\left(k\right)}:=G/N_{k}$ is projective in $\mathbf{LieAb}$ for every $k\in\omega$. Then by Lemma LABEL:Lemma:connected-component we have that $c\left(G\right)=G_{\mathbb{R}}$ and $G\left/c\left(G\right)\right.\cong\mathrm{\mathrm{lim}}_{k}G_{\mathbb{Z}}^{\left(k\right)}$ where $(G_{\mathbb{Z}}^{\left(k\right)})_{k\in\omega}$ is an epimorphic tower of free abelian groups. Whence, $G_{\mathbb{Z}}$ is isomorphic to a product of free abelian groups and hence projective in $\mathbf{proLiePAb}$. Hence, $G\cong G_{\mathbb{R}}\oplus G_{\mathbb{Z}}$ is projective in $\mathbf{proLiePAb}$.
@@
We have that $\mathbf{proLiePAb}$ has enough projectives by Theorem 3.51 and Lemma 3.44. We now prove that closed subgroups of projectives are projective. Suppose that $G$ is a closed subgroup of $\left(F\oplus\mathbb{R}\right)^{\omega}$ where $F$ is a countable free abelian group of infinite rank. Then we have that $G_{\mathbb{S}^{1}}=0$ and $G_{\mathrm{t}}=0$. Hence, we have that $G\cong G_{\mathbb{R}}\oplus G_{\mathbb{Z}}$. Since $G_{\mathbb{R}}$ is injective, we can assume without loss of generality that $G=G_{\mathbb{Z}}$.
@@
We consider the cohomological derived functor of $\mathrm{H}^{0}\circ\mathrm{Hom}^{\bullet}$ for locally compact Polish abelian groups introduced in [5]; see also [22, Section 4]. For every locally compact Polish abelian group $A$, we have an exact sequence
@@
Lemma 3.55.
Every countable divisible group is injective in the category of non-Archimedean Polish abelian groups.
Proof.
Let $D$ be a countable divisible group, and $A$ be a non-Archimedean Polish abelian group. Then there exists a sequence $\left(C_{i}\right)$ of countable abelian groups such that, setting
$C:=\prod_{n}C_{n}\text{,}$
there exists an injective continuous homomorphism $A\rightarrow C$. This induces a surjective homomorphism
$\mathrm{Ext}\left(C,D\right)\rightarrow\mathrm{Ext}\left(A,D\right)\text{.}$
As $D$ is injective in the category of countable abelian groups, for every $i\in\omega$ we have that $\mathrm{Ext}\left(C_{i},D\right)=0$. By Lemma 3.47 this implies that $\mathrm{Ext}\left(C,D\right)=0$ and hence $\mathrm{Ext}\left(A,D\right)=0$. ∎
@@
Thus, we have that $D=D\cap\mathbb{T}^{\omega}\subseteq\mathbb{T}^{\omega}$. Thus, we have that the inclusion $D\rightarrow\mathbb{T}^{\omega}$ induces a surjective homomorphism $\mathbb{Z}^{\left(\omega\right)}\rightarrow D^{\vee}$. Since $D^{\vee}$ is compact and connected, it is uncountable whenever it is nonzero. Hence, $D^{\vee}=0$ and $D=0$. ∎
@@ 4.1. Pro-$p$ Polish abelian groups
Let us say that a pro-$p$ Polish group is a pro-Lie Polish abelian topological $p$-group. We let $\mathbf{PAb}\left(p\right)$ be the category of pro-$p$ Polish abelian groups, which is easily seen to be a thick subcategory of $\mathbf{proLiePAb}$. A locally compact Polish abelian group $G$ is pro-$p$ if and only if it has a basis of zero neighborhoods consisting of subgroups $U$ such that $G/U$ is a $p$-group. A pro-Lie Polish abelian group is a pro-$p$ group if and only if $G/N$ is pro-$p$ for every $N\in\mathcal{N}\left(G\right)$.
@@
For an ordinal $\alpha$, as a particular instance of [33, Theorem 1.5] in the case of the cotorsion functor with enough projectives $G\mapsto p^{\alpha}G$, we have the following:
@@
due to the fact that $G/H$ is countable, it suffices to prove that, for every $k\in\omega$, the image of $[c]$ in $\mathrm{Ext}(\mathbb{Z}/p\mathbb{Z},H)$ induced by the inclusion $\mathbb{Z}/p\mathbb{Z}\rightarrow\left(\mathbb{Z}/p\mathbb{Z}\right)^{\mathbb{N}}$ in the $k$-th coordinate, is nonzero. As this image is the element $[c_{k}]$ represented by the cocycle $c_{k}$, we have that it is nontrivial by the choice of $c_{k}$. This shows that $\mathrm{Ext}(\left(\mathbb{Z}/p\mathbb{Z}\right)^{\mathbb{N}},G)\neq 0$.
@@
(1$\Rightarrow$(2) By Theorem 4.2, it suffices to prove that $G$ is divisible and has a basis of zero neighborhoods consisting of divisible subgroups. As in the proof of (1$\Rightarrow$(2) of Theorem 4.2, we have that $G$ is divisible. Suppose that $G$ does not have a basis of zero neighborhoods consisting of divisible subgroups. Then it has a basis $\left(V_{k}\right)_{k\in\omega}$ of compact open subgroups that are not divisible and such that $d\left(V_{k}\right)$ is not open. For $k\in\omega$, let $\rho_{k}\geq 1$ be the least countable ordinal such that $p^{\rho_{k}}V_{k}$ is not open. By compactness of $V_{k}$, we have that $\rho_{k}\leq\omega$. Indeed, if $\rho_{k}>\omega$, then we have that $\left\{p^{n}V\setminus p^{n+1}V:n\in\omega\right\}$ is an infinite clopen partition of the compact set $V\setminus\rho^{\omega}V$, which is impossible. Thus, proceeding as in the proof of (1$\Rightarrow$(2) of Theorem 4.2, one obtains that there exists a sequence $\left(T_{k}\right)$ of finite $p$-groups such that, setting $T:=\prod_{k\in\omega}T_{k}$, one has that $\mathrm{Ext}\left(T,G\right)\neq 0$, which is a contradiction.
@@
Suppose that $A$ is an injective pro-$p$ Polish abelian group, $B$ is a Polish abelian group, and $\pi:A\rightarrow B$ is a surjective continuous homomorphism. Since $A$ is divisible, $B$ is also divisible. If $\left(V_{k}\right)$ is a basis of zero neighborhoods of $A$ consisting of divisible subgroups, then $\left(\pi\left(V_{k}\right)\right)$ is a basis of zero neighborhoods of $B$ consisting of divisible subgroups by the Open Mapping Theorem for Polish groups. Thus, $B$ is an injective pro-$p$ Polish abelian group. This shows that quotients of injectives are injective in $\mathbf{PAb}\left(p\right)$, and hence $\mathbf{PAb}\left(p\right)$ has homological dimension at most $1$. ∎